40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 29 21:02:45 2012, Jenne, Update, IOO, Optics between Faraday and PRM are centered, realigned 
    Reply  Thu Aug 30 00:08:34 2012, jamie, Update, General, In vac plans for tomorrow, 8/30 
       Reply  Thu Aug 30 08:12:39 2012, Steve, Update, General, In vac plans for tomorrow, 8/30 
Message ID: 7313     Entry time: Wed Aug 29 21:02:45 2012     Reply to this: 7314
Author: Jenne 
Type: Update 
Category: IOO 
Subject: Optics between Faraday and PRM are centered, realigned 

[Jenne, Suresh, with support from Jamie and Koji]

MC spots measured, MC1, MC3 no change.

No clipping going through Faraday.

Beam hitting to the right of center of PZT1.  It was translated sideways so we are now hitting it on the center.  Knobs adjusted so we hit center of MMT1.

Beam totally obscured by Faraday on the way to MMT2.  MMT2 moved north, so that we clear the Faraday by more than a beam diameter.  MMT1 knobs adjusted to hit center of MMT2.

MMT2 knobs adjusted to hit center of PZT2.

PZT2 didn't have enough range with knobs, so we loosened it, pointed then adjusted with knobs so we're hitting center of PRM. 

We need to check spot centering on PRM with camera tomorrow.

Suresh checked that we're not clipped by IP ANG/POS pickoff mirrors, but we haven't done any alignment of IP ANG/POS.

 

Tomorrow:  Open ITMX door.  Check with Watek that we're hitting center of PRM.  Then look to see if we're hitting center of PR2.  Then, continue through the chain of optics.

ELOG V3.1.3-